OpenStudy (muscrat123):

i have a word problem on systems of equations that i am completely lost on. my teacher says i have the right setup, i just need to eliminate a variable, which i know how to do, but im stuck on this question. HELP!!!!!!!!! Will MEDAL AND FAN will post question once i have help Molly's Custom Kitchen Supplies sells homemade forks and spoons. It costs the store $1.70 to buy the supplies to make a fork and $1.30 to buy the supplies to make a spoon. The store sells forks for $5.60 and spoons for 5.40. Last April Molly's Custom Kitchen Supplies spent $37.90 on materials for forks and spoons. They sold the finished products for a total of $147.20. How many forks and how many spoons did they make last April?

imqwerty (imqwerty):

ok :)

imqwerty (imqwerty):

kay :)

imqwerty (imqwerty):

yea m jst helpin some other ppl too :) i'll come back as soon as i get notification :)

imqwerty (imqwerty):

(:

OpenStudy (muscrat123):

mmk. what i have so far is this: 1.70f + 1.30s = $37.90 5.60f + 5.40 = 147.20 my teacher says this is correct, i just need to evaluate now, and im stuck @imqwerty

OpenStudy (muscrat123):

@imqwerty

OpenStudy (muscrat123):

k im back

imqwerty (imqwerty):

ok there is a typo that shuld be 5.40s

OpenStudy (muscrat123):

yeaa sorry i was in a rush but yea thats wat its supposed to say

imqwerty (imqwerty):

np :)

imqwerty (imqwerty):

1.70f + 1.30s = $37.90 1.7f=37.9-1.3s f=(37.9-1.3s)/1.7 substitute this value of f in the second equation and u'll get the value of s then put the value of s in any of these equations to get the value of f

OpenStudy (muscrat123):

idunderstand @imqwerty

imqwerty (imqwerty):

did u get how we took out the value of f in terms of s ?

OpenStudy (muscrat123):

no do u know algebra nation?

imqwerty (imqwerty):

no

OpenStudy (muscrat123):

whelp...we use the elimination method, elimination multiple, and the substitution method to solve this. we havent learned that way

imqwerty (imqwerty):

ok lets do it that way

OpenStudy (muscrat123):

YAY!

OpenStudy (muscrat123):

do u know how?

OpenStudy (muscrat123):

?

imqwerty (imqwerty):

1.70f + 1.30s = $37.90 multiply this equation by 5.6/1.7 u get- (5.6/1.7)1.7f+(5.6/1.7)1.3s=(5.6/1.7)37.9 5.6f+(5.6/1.7)1.3s=(5.6/1.7)37.9 now subtract this equation from the second equation to eliminate 5.6f :) and then u'll be left with an equation with s.. can u solve it now?

OpenStudy (muscrat123):

?

OpenStudy (muscrat123):

imust seem really stupid, i know how to do this, just not for this problem sorry

imqwerty (imqwerty):

don wrry :)

OpenStudy (muscrat123):

can u still help me?

imqwerty (imqwerty):

yea ofc

OpenStudy (muscrat123):

kay thx

imqwerty (imqwerty):

so we have this- 5.6f+(5.6/1.7)1.3s=(5.6/1.7)37.9 equation1 and we got another equation-5.60f + 5.40s = 147.20 equation 2 subtracting equation 1 frm 2 u get- 5.60f + 5.40s - [5.6f+(5.6/1.7)1.3s] =147.20 -(5.6/1.7)37.9 5.60f + 5.40s - 5.6f - (5.6/1.7)1.3s = 147.20 -(5.6/1.7)37.9 5.4s-(5.6/1.7)1.3s =147.2 -(5.6/1.7)37.9 can u solve this equation to get s if ur stuck then don hesitate to ask :P :)

OpenStudy (muscrat123):

why are u dividing?

imqwerty (imqwerty):

u mean that 5.6/1.7?

OpenStudy (muscrat123):

yea

imqwerty (imqwerty):

well we wanted to finish off the 5.6 f in the second equation by we had a 1.7f in the 1st equation so to make it 5.6f we multiplied the whole equation by 5.6/1.7 :)

OpenStudy (muscrat123):

i dont understand

OpenStudy (muscrat123):

i dont hae time, i got a TON of hw

OpenStudy (muscrat123):

i guess ill just leave it blank, i have no idea what ur saying

OpenStudy (muscrat123):

@Hero